if the cost of production of milk per tin increased by 50 k (from $2.00 to $2.50) and the quantity increased from 200 to 280 tins per week, determined the cost of elasticity of supply.
1
Expert's answer
2020-04-24T11:02:20-0400

Elasticity formula:

"Es=\frac{Q2-Q1}{Q2+Q1}:\frac{P2-P1}{P2+P1}."

So,

"Es= \frac{280-200}{280+200}:\frac{2.50-2}{2.5+2}=\frac{1}{6}\times\frac{9}{1}=\frac{3}{2}."